OpenStudy (anonymous):

HELP PLSSSSSSSS The total snowfall in Buffalo, NY one year was one hundred twenty inches. If the total snowfall the following year was ninety inches, what was the percent change? Show and explain your solution.

OpenStudy (kittyjosh):

A percent of change is a ratio that compares the change in quantity to the original amount. The formula for percent of change is: \[Percent of Change = \frac{ Amount of Change }{ Original Amount }\] An example of this would be: Original: 12, New: 9. \[\frac{ 12-9 }{ 12 }=\frac{ 3 }{12 }=\frac{ 1 }{ 4 }\] One Fourth or 25% So, use this equation to solve it. I will help you start... Here is your equation: \[\frac{ 120-90 }{ 120 }=\frac{ 30 }{ 120 }=\frac{ ? }{ ? }\] Good Luck, and Happy Learning! =)

OpenStudy (anonymous):

whats the answer doe? @kittyjosh

OpenStudy (kittyjosh):

Just divide 30 by 120, and convert it to a percentage....

OpenStudy (anonymous):

4 percent? @KittyJosh

OpenStudy (kittyjosh):

30/120 = .25*100 = 25% Decrease

OpenStudy (anonymous):

im confused @KittyJosh

OpenStudy (kittyjosh):

About what?